OpenStudy (anonymous):

What is the value of x? x/4-5=-6

pooja195 (pooja195):

Add 5 to both sides since we are islating x -6+5=?

OpenStudy (anonymous):

-1

pooja195 (pooja195):

x/4=-1 now multiply by the reciprocal x/4 ---> 4/x -4 x 4=?

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

gta5 lol have you heard of vanossgaming?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

Im soo sorry >.< i meant 4 x -1 = ?

OpenStudy (anonymous):

omg THERE ARE MY FAV YOUTUBER

OpenStudy (anonymous):

-4 ;)

pooja195 (pooja195):

-_-

OpenStudy (anonymous):

-4

pooja195 (pooja195):

Grats thats your answer. \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

pooja dont beat yourself up XD

OpenStudy (anonymous):

ty so much i will fan and medal you

pooja195 (pooja195):

No probs :)